OpenStudy (anonymous):

Solve the equation giving your answer correct to 3 significant figures. ln(1 + x2) = 1 + 2 ln x

OpenStudy (anonymous):

\[\ln(1 + x^2) = 1 + 2 \ln x\]

OpenStudy (anonymous):

Numerical analysis a good way ... do you know Newton Raphson?

OpenStudy (anonymous):

Well \( \large x = \frac{1}{\sqrt{-1+e}} \)

OpenStudy (lalaly):

\[\large{e^{\ln(1+x^2)}}=e^{1+2lnx}\]\[\large{1+x^2=e^1e^{2lnx}}\]\[\large{1+x^2=e^1e^{lnx^2}}\]\[\large{1+x^2=e^1x^2}\]\[x^2(e^1-1)=1\]\[x^2=\frac{1}{e^1-1}\]

OpenStudy (anonymous):

Um.. not following

OpenStudy (anonymous):

um.. where are you struck?

OpenStudy (anonymous):

I'm stuck on... well... everything?

OpenStudy (anonymous):

Did you read lala's answer? She explained it quite well.

OpenStudy (anonymous):

Yes, I see it... but can you explain it a bit slower? I'm not very good with the In equations...

OpenStudy (anonymous):

Do you know \( \Huge a^{ \log_a b } = b \)

OpenStudy (anonymous):

Now I do. :)

OpenStudy (anonymous):

that's great, now you can figure out the rest ?

OpenStudy (anonymous):

I'd still like some help if you would?

OpenStudy (anonymous):

Btw the final answer is \( 0.763 \)

OpenStudy (anonymous):

hm sure :)

OpenStudy (anonymous):

Yup, the answer is here... but I just don't know how to arrive at it...

OpenStudy (anonymous):

where are you stuck again?

OpenStudy (anonymous):

the last two parts of lalaly's steps.

OpenStudy (anonymous):

Actually, I understand that

OpenStudy (anonymous):

How do you get the answer?

OpenStudy (anonymous):

I used Mathematica

OpenStudy (anonymous):

Thanks. I got it.

OpenStudy (anonymous):

Glad to help :)

OpenStudy (phi):

These videos might be too boring http://www.khanacademy.org/?video=introduction-to-logarithm-properties#logarithms but they are helpful if you need review, or have someone walk through the properties of logs

OpenStudy (anonymous):

I'll check it out when I have time! :)